### v1

@@ D.2 Pass-Through Heterogeneity
Table D.2 examines the heterogeneity in tariff pass-through through specification (3.2) by including buyer-by-year fixed effects ($FE_{jt}$). This more demanding specification accounts for time-varying shocks at the buyer level, while also controlling for product–year and exporter country–year fixed effects ($FE_{ht}+FE_{ct}+FE_{jt}$). Findings replicate the results of incomplete tariff pass-through reported in Table 3, suggesting that exporters adjust marginal costs in response to demand shifts from dominant importers, thereby absorbing a substantial fraction of tariff shocks. These results demonstrate the critical role of the cost channel as, by Proposition 3, the pass-through decreases with the buyer share $x_{ij}$.
@@
$\Delta\ln(1+\tau_{cht})\cdot\ln\text{longevity}_{ijht}$ 0.086 0.097
@@
$\Delta\ln(1+\tau_{cht})\cdot s_{ijht-1}$ 0.023 0.029 0.018
@@
$\Delta\ln(1+\tau_{cht})\cdot x_{ijht-1}$0.271 -0.271 -0.280
@@
Notes: This table reports estimates of the pass-through of statutory tariffs, $\Delta\ln(1+\tau_{cht})$, to duty-exclusive prices at the exporter–importer–product–year level, $\Delta\ln p_{ijht}$. Columns (2) and (6) interact tariffs with the log of relationship longevity, measured as the number of years that buyer $j$ and supplier $i$ have transacted in product $h$. Columns (3) and (5) interact tariffs with the lagged supplier share, $s_{ijht-1}$, defined as supplier $i$’s share in buyer $j$’s imports of product $h$. Columns (4) and (5) interact tariffs with the lagged buyer share, $x_{ijht-1}$, defined as buyer $j$’s share in supplier $i$’s exports of product $h$. All regressions include product–year, exporter country–year, and importer–year fixed effects ($FE_{ht}+FE_{ct}+FE_{jt}$). Controls include: (i) $\ln\text{longevity}_{ijht}$; (ii) $\Delta\ln q_{i(-j)ht}$, the change in exporter $i$’s total sales of $h$ to U.S. buyers other than $j$; and (iii) $\Delta\ln p_{(-i)jht}$, the weighted average price change charged by other suppliers of $h$ to buyer $j$, using lagged shares as weights. Standard errors are clustered at the HS8 product and exporter-country level. The sample corresponds to the ”+ Supplier Multi-Buyer” definition in Table C.1. Observation counts are rounded to four significant digits per U.S. Census Bureau disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
Nonlinear Effects
@@
$\Delta\ln p_{ijht}=$ $\alpha_{0}+\alpha_{1}\Delta\ln(1+\tau_{cht})+\sum_{q=2}^{Q}\alpha_{s,q}\cdot\Delta\ln(1+\tau_{cht})\cdot\mathbf{1}\{s_{ijh,t-1}\in Q_{q}\}$
@@
The results reveal no evidence of nonlinearities in pass-through with respect to supplier shares (in purple). Interaction coefficients across the upper quartiles of $s_{ijht}$ are uniformly positive, but small in magnitude and statistically insignificant, indicating that supplier concentration does not materially affect the degree of pass-through. In contrast, buyer shares exhibit a strong, monotonic relationship (in green): pass-through declines significantly at higher quartiles of $x_{ijht}$, consistent with the model’s prediction that dominant buyers constrain suppliers’ ability to shift cost shocks.
@@
Additional Robustness: GE Controls and Price Definitions
Table D.3 presents robustness checks using alternative price definitions and specifications. Columns (1)-(2) exclude the general equilibrium controls from our baseline model. Columns (3)-(4) use duty-exclusive prices that include charges, while Columns (5)-(6) use tariff-inclusive prices. In all cases, we interact tariff changes with lagged supplier and buyer shares, and hold fixed effects constant across specifications for comparability. Across all variations, pass-through estimates and their interaction effects with bilateral market shares remain stable in sign and magnitude, supporting the robustness of the main findings.
Table D.3: Additional Robustness: GE Controls and Price Definitions
Dependent variable: $\ln p_{ijht}$ (excl. GE controls) $\ln p^{c}_{ijht}$ (before duty, incl. charges) $\ln p^{\text{duty}}_{ijht}$ (tariff inclusive)
@@
$\Delta\ln(1+\tau_{cht})\cdot s_{ijht-1}$ 0.048 0.036 0.033 0.004 0.017 0.007
@@
$\Delta\ln(1+\tau_{cht})\cdot x_{ijht-1}$0.399 -0.278 -0.407 -0.280 -0.421 -0.287
@@
Notes: This table reports robustness checks on tariff pass-through specifications using alternative price definitions and control sets. Columns (1)–(2) exclude general equilibrium controls; Columns (3)–(4) use pre-duty prices including charges; and Columns (5)–(6) use tariff-inclusive prices. In each case, we report specifications using either baseline fixed effects ($FE_{ht}+FE_{cs}$) or a more stringent set of fixed effects ($FE_{ht}+FE_{ct}+FE_{jt}$). Standard errors are clustered at the HS8 product and exporter-country level. The number of observations is rounded to four significant digits in accordance with U.S. Census Bureau disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
@@ D.3 Additional Results on Model Estimation
We assess the robustness of our structural estimates in Section 4.2. We first replicate the main GMM estimation using an expanded sample that includes all Broad Economic Categories (BECs), rather than excluding consumption goods as in the baseline. This extended product scope allows us to test whether our key parameter estimates, namely, returns to scale ($\theta$) and relative bargaining power ($\phi$), are sensitive to the exclusion of consumer-oriented products. As shown in Table D.4, the results remain fairly stable, suggesting that the baseline findings are not driven by product composition.
We then examine how sensitive the estimates are to alternative values of calibrated model parameters. First, we vary the elasticity of substitution across foreign varieties ($\rho$) by setting $\rho=5$ instead of 10, which is consistent with the lower end of the estimates in the literature. Second, we relax the assumption of constant returns to scale for the importers and set $\varrho=0.5$ instead of 1. We report the results in Table D.5. Column (1) shows the estimated values when setting $\rho=5$, and Column (2) shows the estimated values when setting $\varrho=0.5$. Furthermore, in Table D.5, we also estimate the parameters using an alternative sample constructed by utilizing the RPT indicator from LFTTD. [^37] See Section 3.3 for the discussion on selection through RPT indicators. Columns (3) and (4) report the estimated values for this set of sample. Throughout these alternative setups,the resulting estimates remain robust, suggesting that the estimated values are not sensitive to particular values of other parameters or set of sample.
@@
Panel A: Calibrated Parameters $\hat{\nu}$ $\hat{\gamma}$ $\hat{\rho}$ 4 0.5 10 Panel B: Estimated Parameters (GMM) (1) (2) (3) (4) Rel. bargaining power: $\ln\widehat{\frac{\phi}{1-\phi}}$ 1.162 0.558 (0.026) (0.022) Returns to scale ($\hat{\theta}$) 0.505 0.573 0.432 0.586 (0.003) (0.005) (0.004) (0.005) Constant 2.780 0.742 (0.190) (0.083) Longevity 0.061 0.742 (0.023) (0.069) Number of HS10 transactions -0.240 -0.019 (0.019) (0.011) Multiple HS10 dummy 0.080 0.190 (0.026) (0.025) Lagged outside option -0.188 -0.237 (0.017) (0.021) None Yes Yes No No $FE_{h}+FE_{t}+FE_{j}$ No No Yes Yes Observations 6,143,000 Panel C: Implied Bargaining Powers ($\hat{\phi}$) Mean 0.762 0.896 0.636 0.848 (0.005) (0.068) (0.005) (0.112) Median 0.913 0.878 (0.068) (0.112) Notes: This table presents model estimates based on a sample that includes all Broad Economic Categories (BEC), including consumption goods, for the period 2001-2016. Panel A reports calibrated parameters: the elasticity of demand ($\nu$), the elasticity of costs with respect to foreign input prices ($\gamma$), and the elasticity of substitution across foreign varieties ($\rho$). We set $\varrho=1$, so that $\eta=2.5$. Panel B presents GMM estimates. Columns (1) and (3) impose a constant $\phi$ across bilateral pairs, while Columns (2) and (4) estimate the full vector $\boldsymbol{\kappa}$ to allow for heterogeneity in bargaining power. Specifications differ in the inclusion of fixed effects. Controls include: (i) the log of relationship longevity between exporter $i$ and importer $j$; (ii) the log of the number of transactions between $i$ and $j$ in a given year; (iii) the log of the relative outside option, defined as the ratio of exporter $i$’s sales to other U.S. buyers (excluding $j$) over importer $j$’s purchases from other suppliers (excluding $i$), both in year $t{-}1$; and (iv) a dummy variable equal to one if the $i–j$ pair transacts in more than one HS10 product. Panel C reports the mean and median of the implied bargaining power. Standard errors are robust; those in Panel C are computed using the delta method. The set of instruments includes the number of exporters and importers at the HS10 level, as well as lagged bilateral shares (excluding the focal pair). The number of observations is rounded to four significant digits in accordance with U.S. Census Bureau disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
Table D.5: Robustness of Model Estimates
@@
$\rho=5$ $\varrho=0.5$ Related Party Trade LFTTD (1) (2) (3) (4) Rel. bargaining power: $\ln\widehat{\frac{\phi}{1-\phi}}$ 1.455 1.838 1.618 0.892 (0.038) (0.052) (0.062) (0.047) Returns to scale ($\theta$) 0.427 0.473 0.453 0.381 (0.004) (0.003) (0.005) (0.007) Mean $\hat{\phi}$ 0.811 0.863 0.835 0.709 (0.006) (0.006) (0.009) (0.010) Median $\hat{\phi}$ 0.811 0.863 0.835 0.709 (0.000) (0.000) (0.000) (0.000) None Yes Yes Yes No $FE_{h}+FE_{t}+FE_{j}$ No No No Yes Observations 3,120,000
Notes: This table reports robustness checks for the main model estimates. The columns explore sensitivity to changes in key calibrated parameters and sample definitions. Column (1) varies the elasticity of substitution across foreign varieties ($\rho$); Column (2) changes the downstream returns to scale parameter ($\varrho$). Columns (3)–(4) use a sample based on related-party indicators from LFTTD (RPT), without and with fixed effects for buyer, product, and time. The number of observations is rounded to four significant digits in accordance with U.S. Census Bureau disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
@@ D.4 Additional Results on Model Fit
Sensitivity to alternative parameter values.
Table D.6 presents IV-based goodness-of-fit tests as in equation (4.5) for an alternative set of calibrated parameters of the model and two alternative sets of fixed effects. The results show that models incorporating bargaining and decreasing returns to scale (Columns (1) and (5)) have coefficients closer to one, indicating empirical alignment, consistent with the results in Figure 3.
Quantity responses and relationship heterogeneity: data vs. model.
@@
Testing the model-predicted quantity changes.
Panel A in Table D.8 presents IV-based tests comparing observed quantity changes to those predicted by the model under alternative parameterizations. The model-predicted change in quantity is given by:
@@
where $\varepsilon_{ijht}$ is the match-specific demand elasticity and $\widehat{\Delta\ln q_{ijht}}$ is as in equation (4.5). Although formal tests reject all models, the baseline model with bargaining and decreasing returns (Columns (1) and (2)) demonstrates the strongest fit, indicating that this specification best captures the underlying mechanisms of tariff-induced quantity adjustments.
Testing the model-predicted sales changes.
Panel B in Table D.8 provides analogous IV-based tests for observed sales changes computed as:
@@
The baseline specification (Columns (1) and (2)) shows superior alignment between predicted and observed sales, reinforcing the conclusion that the bargaining model with decreasing returns most effectively matches empirical patterns in the data.
@@
$\widehat{\Delta\ln p_{ijht}}$ 1.211 0.781 0.735 1.368 1.247 0.871 0.740 1.544
@@
$\widehat{\Delta\ln p_{ijht}}$ 1.000 0.648 0.612 1.132 1.047 0.720 0.615 1.302
@@
Notes: Each column reports the coefficient from an IV regression of the observed change in log price on the corresponding model-predicted change $\widehat{\Delta\ln p_{ijht}}$, using statutory tariff changes as instruments. Columns (1)-(4) use predicted prices using a $\rho=5$ instead of $\rho=10$ as in the baseline. Columns (5)-(8) use predicted prices using a $\varrho=0.5$ instead of $\varrho=1$ as in the baseline. Notice that in each case, we re-estimate the calibrated parameters $\phi$ and $\theta$, reported in Appendix D.3 In Panel A, all columns include product-time and country-sector fixed effects ($FE_{ht}+FE_{cs}$). In Panel B, all columns include product-time, country-time, and buyer-time fixed effects ($FE_{ht}+FE_{ct}+FE_{jt}$). Standard errors are clustered at the product and exporter-country level. The number of observations is rounded to four significant digits in accordance with U.S. Census Bureau disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
@@
Panel A: Data
@@
$\Delta\ln(1+\tau_{cht})\cdot s_{ijht-1}$0.511 -0.278
(0.323) (0.276)
$\Delta\ln(1+\tau_{cht})\cdot x_{ijht-1}$0.201 -0.130
(0.229) (0.242)
R-squared 0.06 0.10 0.32 0.36
Panel B: Model
(1) (2) (3) (4)
$\Delta\ln(1+\tau_{cht})$1.962 -3.967 -2.169 -3.984

### v2

@@ D.2 Pass-Through Heterogeneity
Table D.2 examines heterogeneity in tariff pass-through using specification (3.2) with a more demanding set of fixed effects. In addition to product–time fixed effects ($FE_{ht}$), the specification includes importer–time and exporting country–time fixed effects ($FE_{jt}+FE_{ct}$), accounting for time-varying shocks at the buyer level while flexibly controlling for exporter-side conditions.
Findings replicate the incomplete tariff pass-through estimates reported in Table 3, consistent with exporters adjusting marginal costs in response to demand shifts from large buyers and thereby absorbing a substantial fraction of tariff shocks. In line with Proposition 3, pass-through declines with the buyer share $x_{ij}$, highlighting the central role of the cost channel.
@@
$\Delta\ln(1+\tau_{cht})\cdot\ln\text{longevity}_{ijht}$ 0.088 0.101
@@
$\Delta\ln(1+\tau_{cht})\cdot s_{ijht-1}$ -0.003 0.004 -0.007
@@
$\Delta\ln(1+\tau_{cht})\cdot x_{ijht-1}$ -0.282 -0.280 -0.289
@@
Notes: This table reports estimates of the pass-through of statutory tariffs, $\Delta\ln(1+\tau_{cht})$, to duty-exclusive CIF prices at the exporter–importer–product–year level, $\Delta\ln p_{ijht}$. Columns (2) and (6) interact tariffs with the log of relationship longevity, measured as the number of years that buyer $j$ and supplier $i$ have transacted in product $h$. Columns (3) and (5) interact tariffs with the lagged supplier share, $s_{ijht-1}$, defined as supplier $i$’s share in buyer $j$’s imports of product $h$. Columns (4) and (5) interact tariffs with the lagged buyer share, $x_{ijht-1}$, defined as buyer $j$’s share in supplier $i$’s exports of product $h$. All regressions include product–year, exporter country–year, and importer–year fixed effects ($FE_{ht}+FE_{ct}+FE_{jt}$). Controls include: (i) $\ln\text{longevity}_{ijht}$; (ii) $\Delta\ln q_{i(-j)ht}$, the change in exporter $i$’s total sales of $h$ to U.S. buyers other than $j$; and (iii) $\Delta\ln p_{(-i)jht}$, the weighted average price change charged by other suppliers of $h$ to buyer $j$, using lagged shares as weights. Standard errors are clustered at the HS8 product and exporter-country level. The sample corresponds to the ”+ Supplier Multi-Buyer” definition in Table C.1. Observation counts are rounded to four significant digits per U.S. Census Bureau disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
Nonlinear Effects.
@@
$\Delta\ln p_{ijht}=\$$\alpha_{0}+\alpha_{1}\Delta\ln(1+\tau_{cht})+\sum_{q=2}^{Q}\alpha_{s,q}\cdot\Delta\ln(1+\tau_{cht})\cdot\mathbf{1}\{s_{ijh,t-1}\in Q_{q}\}$
@@
The results reveal no evidence of nonlinearities in pass-through with respect to supplier shares (in purple). Interaction coefficients across the upper quartiles of $s_{ijht}$ are uniformly positive, but small and statistically insignificant, indicating that supplier concentration does not materially affect the degree of pass-through. In contrast, buyer shares display a strong monotonic pattern (in green): pass-through declines significantly across higher quartiles of $x_{ijht}$, consistent with the model’s prediction that dominant buyers limit suppliers’ ability to transmit cost shocks.
@@
Additional Robustness: GE Controls and Price Definitions.
Table D.3 presents robustness checks using alternative price definitions and specifications. Columns (1)-(2) exclude the general equilibrium controls included in the baseline model. Columns (3)-(4) use duty-exclusive prices that exclude charges, while Columns (5)-(6) use tariff-inclusive prices. In all cases, tariff changes are interacted with lagged supplier and buyer shares, and the set of fixed effects is held constant to ensure comparability across specifications. Estimates remain stable in sign and magnitude across all variants, confirming the robustness of the main results.
Table D.3: Additional Robustness: General Equilibrium Controls and Price Definitions
Dependent variable: $\ln p_{ijht}$ (excl. GE controls) $\ln p^{c}_{ijht}$ (before duty, excl. charges) $\ln p^{\text{duty}}_{ijht}$ (tariff inclusive)
@@
$\Delta\ln(1+\tau_{cht})\cdot s_{ijht-1}$ 0.027 0.011 0.054 0.029 0.017 0.007
@@
$\Delta\ln(1+\tau_{cht})\cdot x_{ijht-1}$ -0.403 -0.287 -0.403 -0.271 -0.421 -0.287
@@
Notes: This table reports robustness checks on tariff pass-through specifications using alternative price definitions and control sets. Columns (1)–(2) exclude general equilibrium controls; Columns (3)–(4) use pre-duty prices excluding charges; and Columns (5)–(6) use tariff-inclusive prices. In each case, we report specifications using either baseline fixed effects ($FE_{ht}+FE_{cs}$) or a more stringent set of fixed effects ($FE_{ht}+FE_{ct}+FE_{jt}$). Standard errors are clustered at the HS8 product and exporter-country level. The number of observations is rounded to four significant digits in accordance with U.S. Census Bureau disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
@@ D.3 Additional Results on Model Estimation
We assess the robustness of our structural estimates in Section 4.2. We first re-estimate the main GMM specification using an expanded sample that includes all Broad Economic Categories (BECs). We also estimate the structural parameters on a sample restricted to consumption goods only. Results are reported in Tables D.4 and D.5, respectively
Estimated returns to scale are systematically higher for consumption goods. Restricting the estimation to consumption goods raises $\hat{\theta}$ from 0.42–0.51 in the baseline to 0.54–0.66. Expanding the sample to include all BEC categories also increases estimates, with $\hat{\theta}$ ranging from 0.48 to 0.60, consistent with the higher values observed for consumption goods. This pattern suggests that consumption goods are less relationship-specific and therefore face more elastic supply conditions, implying greater substitutability across suppliers and higher estimated returns to scale in the model.
Finally, Table D.6 examines the robustness of our structural estimates to an alternative definition of related-party trade. Instead of using our baseline ORBIS-based classification, we re-estimate the model excluding transactions identified as related-party trade using the standard RPT indicator in the U.S. Census data. Appendix C.1 explains the difference between the two indicators. Reassuringly, the estimated bargaining-power and returns-to-scale parameters remain very close to the baseline results, indicating that our main conclusions are not driven by the particular way related-party relationships are classified.
@@
Panel A: Calibrated Parameters $\hat{\nu}$ $\hat{\gamma}$ $\hat{\rho}$ 4 0.5 10 Panel B: Estimated Parameters (GMM) (1) (2) (3) (4) Rel. bargaining power: $\ln\widehat{\frac{\phi}{1-\phi}}$ 1.180 0.940 (0.023) (0.021) Returns to scale ($\hat{\theta}$) 0.508 0.585 0.483 0.595 (0.002) (0.003) (0.003) (0.003) Constant 3.355 1.299 (0.196) (0.103) Longevity 0.115 0.902 (0.026) (0.090) Number of HS10 transactions -0.315 -0.090 (0.021) (0.016) Multiple HS10 dummy 0.137 0.199 (0.028) (0.029) Lagged outside option -0.267 -0.272 (0.019) (0.023) None Yes Yes No No $FE_{h}+FE_{t}+FE_{j}$ No No Yes Yes Observations 6,143,000 Panel C: Implied Bargaining Powers ($\hat{\phi}$) Mean 0.765 0.923 0.719 0.899 (0.004) (0.075) (0.004) (0.095) Median 0.946 0.928 (0.075) (0.095)
@@
Notes: This table presents model estimates based on a sample that includes all Broad Economic Categories (BEC), including consumption goods, for the period 2001-2016. Panel A reports calibrated parameters: demand elasticity ($\nu$), cost elasticity to foreign input prices ($\gamma$), and the elasticity of substitution across foreign varieties ($\rho$). Panel B reports GMM estimates. Columns (1) and (3) impose a constant $\phi$, while Columns (2) and (4) allow heterogeneous bargaining power via $\boldsymbol{\kappa}$. Specifications differ in fixed effects. Controls include relationship longevity, transaction intensity, the relative outside option (lagged), and a multi-product indicator. Panel C reports the mean and median implied bargaining power. Robust standard errors; Panel C uses the delta method. Instruments include HS10-level counts of exporters and importers and lagged bilateral shares (excluding the focal pair). All estimations use CIF duty-exclusive unit values. The number of observations is rounded to four significant digits in accordance with U.S. Census Bureau disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
Table D.5: Estimated Model Primitives - BEC Consumption Categories
Panel A: Calibrated Parameters $\hat{\nu}$ $\hat{\gamma}$ $\hat{\rho}$ 4 0.5 10 Panel B: Estimated Parameters (GMM) (1) (2) (3) (4) Rel. bargaining power: $\ln\widehat{\frac{\phi}{1-\phi}}$ 0.936 0.729 (0.023) (0.021) Returns to scale ($\hat{\theta}$) 0.558 0.646 0.539 0.664 (0.003) (0.004) (0.003) (0.004) Constant 2.501 0.568 (0.213) (0.093) Longevity 0.628 1.414 (0.070) (0.198) Number of HS10 transactions -0.386 -0.137 (0.037) (0.030) Multiple HS10 dummy 0.452 0.344 (0.056) (0.052) Lagged outside option -0.266 -0.299 (0.027) (0.038) None Yes Yes No No $FE_{h}+FE_{t}+FE_{j}$ No No Yes Yes Observations 3,024,000 Panel C: Implied Bargaining Powers ($\hat{\phi}$) Mean 0.718 0.919 0.674 0.891 (0.005) (0.077) (0.005) (0.111) Median 0.942 0.928 (0.077) (0.111)
Notes: This table presents model estimates based on a sample that only includes consumption goods based on Broad Economic Categories (BEC), for the period 2001-2016. Panel A reports calibrated parameters: demand elasticity ($\nu$), cost elasticity to foreign input prices ($\gamma$), and the elasticity of substitution across foreign varieties ($\rho$). Panel B reports GMM estimates. Columns (1) and (3) impose a constant $\phi$, while Columns (2) and (4) allow heterogeneous bargaining power via $\boldsymbol{\kappa}$. Specifications differ in fixed effects. Controls include relationship longevity, transaction intensity, the relative outside option (lagged), and a multi-product indicator. Panel C reports the mean and median implied bargaining power. Robust standard errors; Panel C uses the delta method. Instruments include HS10-level counts of exporters and importers and lagged bilateral shares (excluding the focal pair). All estimations use CIF duty-exclusive unit values. The number of observations is rounded to four significant digits in accordance with U.S. Census Bureau disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
Table D.6: Estimated Model Primitives under Alternative Related-Party Classification
@@
(1) (2) (3) (4) Implied Mean Bargaining Power ($\hat{\phi}$) 0.837 0.928 0.761 0.899 (0.008) (0.078) (0.008) (0.082) Returns to scale ($\hat{\theta}$) 0.455 0.500 0.410 0.504 (0.004) (0.005) (0.005) (0.006) None Yes Yes No No $FE_{h}+FE_{t}+FE_{j}$ No No Yes Yes
Notes: Model estimates are based on U.S. arm’s-length imports of intermediate inputs and capital goods over 2001–2016, excluding transactions identified as related-party trade using the RPT indicator in the U.S. Census data; see Appendix C.1 for details. Columns (1) and (3) impose a constant $\phi$, while Columns (2) and (4) allow for heterogeneous bargaining power through $\boldsymbol{\kappa}$, parameterized as a function of relationship longevity, transaction intensity, the lagged relative outside option, and a multi-product indicator. The demand elasticity is set to $\nu=4$, the cost elasticity to foreign input prices to $\gamma=0.5$, and the elasticity of substitution across foreign varieties to $\rho=10$. Specifications differ in the fixed effects included. Robust standard errors are reported in parentheses. All estimations use CIF duty-exclusive unit values. All regressions include 3,120,000 observations. The number of observations is rounded to four significant digits in accordance with U.S. Census Bureau disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
@@ D.4 Additional Results on Model Fit
Pass-through and Relationship Heterogeneity: Data vs. Model.
Table D.7 reports tariff pass-through in the data and in the model (baseline and restricted variants) under a more demanding set of fixed effects. Panel A shows that the model captures the main pass-through effect in the data very closely. Panel B shows how pass-through varies with buyer and supplier concentration, measured by lagged buyer and supplier shares. The baseline model closely replicates these patterns: pass-through declines sharply with buyer concentration ($x_{ijht-1}$) and increases modestly with supplier concentration ($s_{ijht-1}$), with magnitudes comparable to those observed in the data.
IV-Based Goodness-of-Fit Test.
We formally assess model fit using an IV regression following adao2023putting:
$\Delta\ln p_{ijht}=\beta\,\widehat{\Delta\ln p_{ijht}}+\mathbf{FE}+u_{ijht},$ (D.2)
where $\Delta\ln p_{ijht}$ is the observed change in the duty-inclusive price and $\widehat{\Delta\ln p_{ijht}}$ is the corresponding model-implied change. Because observed price changes may reflect shocks unrelated to tariffs, we instrument $\widehat{\Delta\ln p_{ijht}}$ with statutory tariff changes to isolate tariff-driven variation. [^36] Unlike the pass-through analysis in the main text, which uses duty-exclusive prices to mitigate measurement error in applied duties, the IV strategy isolates variation induced by statutory tariffs and evaluates the model against duty-inclusive price changes. Under the null that the model correctly captures tariff-induced price movements, the IV coefficient $\hat{\beta}$ equals one.
Figure D.2 reports the IV estimates of $\beta$ across the four model variants under the two fixed-effects specifications. Blue markers correspond to regressions with product–time and country–sector fixed effects, while red markers use product–time, country–time, and buyer–time fixed effects. The model generates price changes closely aligned with those observed in the data. Formally, however, we cannot reject any of the variants at conventional levels of significance.
Sensitivity of Model Fit to Alternative Parameter Values.
In our baseline estimation we set $\varrho=1$, $\nu=4$, and $\gamma=0.5$. As all these parameters enter the model only though $\eta$, fixing the buyer’s returns to scale parameter $\varrho$ to one does not restrict generality. Nevertheless, we conduct a robustness exercise where we set $\varrho=0.5$, close to the returns to scale parameter we estimate for the supplier side. Further, we also consider a value of $\rho=5$, the lower end of the range provided in Anderson2004.
Table D.8 presents IV-based goodness-of-fit tests as in equation (D.2) for these alternative set of calibrated parameters of the model and two alternative sets of fixed effects. The results show that models incorporating bargaining and decreasing returns to scale (Columns (1) and (5)) have coefficients closer to one, indicating empirical alignment, consistent with the results in Figure D.2.
Model Fit: Quantity and Sales Responses.
@@
Table D.10 reports IV-based tests comparing observed quantity (Panel A) and sales (Panel B) changes to those implied by the model under alternative parameterizations. The model predicts
@@
where $\varepsilon_{ijht}$ denotes the match-specific demand elasticity and $\widehat{\Delta\ln p_{ijht}}$ is defined in equation (D.2). For sales changes, we use:
@@
Although the formal tests reject all specifications, the specification that combines bargaining with decreasing returns and the specification with decreasing returns only provide the best quantitative fit, consistent with the results for prices.
Table D.7: Price Responses and Relationship Heterogeneity: Data vs. Model
Panel A: Main Effect (Stringent Fixed Effects)
Data Baseline $\phi=0$, $\theta=1$ $\theta=1$ $\phi=0$
$\Delta\ln(1+\tau_{cht})$ -0.184 -0.226 -0.095 -0.014 -0.286
(0.107) (0.009) (0.008) (0.001) (0.011)
R-squared 0.31 0.48 0.42 0.41 0.51
Panel B: Including Interactions (Stringent Fixed Effects)
Data Baseline $\phi=0$, $\theta=1$ $\theta=1$ $\phi=0$
(1) (2) (3) (4) (5)
$\Delta\ln(1+\tau_{cht})$ -0.114 -0.137 -0.030 -0.002 -0.200
(0.104) (0.011) (0.007) (0.001) (0.014)
$\Delta\ln(1+\tau_{cht})\cdot s_{ijht-1}$ 0.004 0.116 -0.120 -0.022 0.073
(0.165) (0.013) (0.022) (0.003) (0.013)
$\Delta\ln(1+\tau_{cht})\cdot x_{ijht-1}$ -0.280 -0.433 -0.078 -0.013 -0.372
(0.134) (0.027) (0.020) (0.004) (0.025)
R-squared 0.31 0.57 0.45 0.45 0.56
$FE_{ht}+FE_{ct}+FE_{jt}$ Yes Yes Yes Yes Yes
Observations 249,000
Notes: This table reports the response of duty-exclusive CIF prices to tariff changes at the exporter–importer–product level. Column (1) presents reduced-form estimates using observed price changes in the data. Columns (2)–(5) report the corresponding responses implied by model-predicted price changes under each specification. Panel A reports the main effect of tariff changes, while Panel B allows the response to vary with lagged supplier shares ($s_{ijht-1}$) and buyer shares ($x_{ijht-1}$). All specifications include product–time, country–time, and buyer-time fixed effects ($FE_{ht}+FE_{ct}+FE_{jt}$). Standard errors are clustered at the HS8 product and exporter-country level. Observation counts are rounded to four significant digits per U.S. Census Bureau disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
Figure D.2: IV-Based Goodness-of-Fit Test
Notes: Each point reports the coefficient from an IV regression of observed log duty-inclusive price changes on model-predicted changes $\widehat{\Delta\ln p_{ijht}}$, using statutory tariffs as instruments. Lines show 95% confidence intervals. Blue and red denote regressions with $\mathrm{FE}_{ht}+\mathrm{FE}_{cs}$ and $\mathrm{FE}_{ht}+\mathrm{FE}_{ct}+\mathrm{FE}_{jt}$, respectively. Standard errors are clustered by product and exporter-country. Observation counts (249,000) are rounded per Census disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
@@
$\widehat{\Delta\ln p_{ijht}}$ 1.158 0.889 0.771 1.391 1.291 0.936 0.776 1.619
@@
$\widehat{\Delta\ln p_{ijht}}$ 0.998 0.776 0.667 1.215 1.128 0.805 0.670 1.432
@@
Notes: Each column reports the coefficient from an IV regression of the observed change in log duty-inclusive price on the corresponding model-predicted change $\widehat{\Delta\ln p_{ijht}}$, using statutory tariff changes as instruments. Columns (1)-(4) use predicted prices using a $\rho=5$ instead of $\rho=10$ as in the baseline. Columns (5)-(8) use predicted prices using a $\varrho=0.5$ instead of $\varrho=1$ as in the baseline. Notice that in each case, we re-estimate the calibrated parameters $\phi$ and $\theta$, reported in Appendix D.3 In Panel A, all columns include product-time and country-sector fixed effects ($FE_{ht}+FE_{cs}$). In Panel B, all columns include product-time, country-time, and buyer-time fixed effects ($FE_{ht}+FE_{ct}+FE_{jt}$). Standard errors are clustered at the product and exporter-country level. The number of observations is rounded to four significant digits in accordance with U.S. Census Bureau disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
@@
Baseline FE Stringent FE
Data Model Data Model
@@
R-squared 0.06 0.25 0.32 0.40
Panel B: Including Interactions
$\Delta\ln(1+\tau_{cht})$ 0.006 -3.683 -0.761 -3.705